2. If D1 = $1.25, g (which is constant) = 4.7%, and P0 = $26.00, what is the stock’s expected dividend yield for the coming year? What is the expected total return for the coming year?
Answer #1

According to Constant Growth DDM model,

P0 = D1/(Re-g)

26 = 1.25/(Re-0.047)

Re-0.047 = 0.0481

Re = 0.0951

Re = 9.51%

..

The expected total return for the coming year = Expected required rate of return

= 9.51%

..

The stock’s expected dividend yield for the coming year = 1.25/26 *100

= 4.81%

Capital Appreciation (capital Gains) = Growth rate

= 4.7%

Total Return = Dividend Yeild + Capital Gains yield

= 4.81+4.7

= 9.51%